PartnerObserving Log for 2007 Oct 09-10

Observers: V. Testa, J. Hill, Stephane Herbert-Fort
Telescope Operator: David Gonzalez Huerta
Telescope and Instrument Support (alseep on-site)

- 25 bias frames

Raining at sunset. Already today we've had problems with LBC Archive disk filling up and with /home on CMU filling up.

Humidity too high until 09:30 UT.

09:30 - Open dome, mirror up only at 95%. Restarted but still problems with
the hardpoints. Try working with the mirror at 95%.

Pointed at N/E (4:30 +35)

09:40 - Built a test OB from the interface and run.
TCS connection error. OB not executed.
Disconnected and reconnected LBT

09:45 - OB started

09:47 - Camera in 'waiting' status
Stopped OB
Turned off and on again
Meanwhile some work on the mirror to try to have it at 100%.

10:10 - Humidity rising, closed again

-- Main.JohnHill - 09 Oct 2007
Topic revision: r4 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback